# Chinese Large Language Models via Unconstrained Generation thanks: * These authors contribute equally

### v2

@@ Abstract
Large language models (LLMs) produce hallucinated text, compromising their practical utility in professional contexts. To assess the reliability of LLMs, numerous initiatives have developed benchmark evaluations for hallucination phenomena. However, they often employ constrained generation techniques to produce the evaluation dataset due to cost and time limitations. For instance, this may involve employing directed hallucination induction or deliberately modifying authentic text to generate hallucinations. These are not congruent with the unrestricted text generation demanded by real-world applications. Furthermore, a well-established Chinese-language dataset dedicated to the evaluation of hallucinations is presently lacking. Consequently, we have developed an $\mathbb{U}$nconstrained $\mathbb{H}$allucination $\mathbb{G}$eneration Evaluation ($\mathbb{UHG}$Eval) benchmark, containing hallucinations generated by LLMs with minimal restrictions. Concurrently, we have established a comprehensive benchmark evaluation framework to aid subsequent researchers in undertaking scalable and reproducible experiments. We have also evaluated prominent Chinese LLMs and the GPT series models to derive insights regarding hallucination.
@@ 1 Introduction
Large language models (LLMs) have unparalleled proficiency in language generation, knowledge application, and intricate reasoning Zhao et al. (2023). However, these models invariably manifest hallucination Rawte et al. (2023), as they often generate content that is incongruent with user input, the model’s output context, or factual information. Real-world hallucination examples from our $\mathbb{UHG}$Eval dataset can be observed in Fig. 1.
@@
While there have been a bunch of efforts to develop benchmarks for hallucination assessment, they always employ restricted techniques to produce particular kinds of hallucinated utterances. This approach is at odds with real-world scenarios where hallucinations arise in unrestricted, spontaneously generated content. For example, HaluEval specifies the type of hallucination in the prompt when generating hallucinated text: “You are trying to answer a question but misunderstand the question context and intention” Li et al. (2023). Additionally, benchmarks such as HADES annotate hallucinations at a finer granularity by generating token-level hallucinations based on text perturbations Liu et al. (2022), but the text perturbation method is still constrained. Besides, many benchmarks are centered on the evaluation of hallucinations in English, neglecting the assessment of such phenomena in Chinese. The extensive lexicon of Chinese characters, combined with the complexities introduced by Chinese word segmentation, renders the Chinese hallucination evaluation particularly arduous and deserving of focused scrutiny.
@@
Our contributions: (1) The development of an unconstrained hallucination evaluation dataset, comprising over 5000 items. Existing methods for constructing datasets often yield biases towards predefined directions, thereby hindering the full simulation of real-world hallucinations. (2) The establishment of a unified and diverse evaluation framework, $\mathbb{UHG}$Eval, that encompasses discriminative, selective, and generative evaluations. Current benchmark methods for hallucination evaluation often exhibit a singular approach and lack task specificity. (3) A comprehensive empirical analysis. We evaluated eight prominent Chinese LLMs and three classic GPT series models to explore the credibility of various LLMs. The aforementioned dataset, evaluation framework, and empirical results collectively constitute the $\mathbb{UHG}$Eval benchmark, which is openly available on GitHub111 Main: https://github.com/IAAR-Shanghai/UHGEval
Dataset creation: https://github.com/IAAR-Shanghai/UHGEval-dataset
2 The UHGEval Dataset
2.1 Data Collection and Pre-processing
Type Share Categories
DOC
27.52%
Politics, Law, Military, Education
NUM
43.34%
Sports, Economy, Market
KNO
6.55%
Science, Technology, Healthcare
GEN
22.59%
Society, Culture, Arts, Entertainment, Weather, Environmental Protection, Disasters, Accidents
2.2 Unconstrained Hallucination Generation
Furthermore, current benchmarks for evaluating hallucination have predominantly relied on a single LLM to produce a hallucinated dataset. Notable examples include HaluEval Li et al. (2023) and PHD Yang et al. (2023b), which exclusively utilize ChatGPT, and FActScore Min et al. (2023) and FACTOR Muhlgay et al. (2023), which solely employ InstructGPT Ouyang et al. (2022). In contrast, our methodology incorporates a suite of five distinct Chinese LLMs to generate hallucinated content. These models include ChatGLM2-6B Du et al. (2022), Baichuan2-13B Yang et al. (2023a), Qwen-14B Bai et al. (2023), InternLM-20B InternLM (2023), and Xinyu-7B. For additional information about the Xinyu series models, please refer to the Appendix B.1.
For each input news article, we concurrently generate five candidate continuations using 5 different LLMs without constraint. Overall, our approach engenders a more unconstrained and heterogeneous generation of hallucinations, mitigating the bias that may arise from the use of a single model or constrained prompting.
2.3 Hallucination Ranking
Algorithm 1 Hallucination ranking
$candidate:list[str]$
$final:str$
$candidate.sort(descend,by=fluency)$
$picked\leftarrow candidate[:3]$triangleright More fluency
$picked.sort(descend,by=likelihood)$
$final\leftarrow picked[0]$triangleright More Hallucination
Hallucination ranking is a crucial step in selecting the most appropriate continuation from a set of candidate continuations generated by LLMs. We employ a ranking process detailed in Algorithm 1. This process relies on two critical metrics: $fluency$, ensuring that the continuation does not become too nonsensical, and $likelihood$, which stands for the likelihood of hallucination occurrence, ensuring that the continuation includes a detectable level of hallucination. By employing such a ranking, it is guaranteed that, in the worst-case scenario, the $final$ candidate ranks at least third in fluency and third in the likelihood of hallucination occurrence, achieving a balanced level. The two metrics are computed as follows.
Fluency
This refers to the coherence and readability of the text. A fluent text should read smoothly, be grammatically correct, and make logical sense in the context of the continuation. To assess fluency, a reward model developed by the Institute for Advanced Algorithms Research (IAAR) is employed, trained to score text quality based on fluency.
Likelihood of Hallucination Occurrence
This metric uses an LLM (e.g., GPT3.5-Turbo) to extract keywords from the continuation and determine whether these keywords have a match in the reference information. The ratio of all matches to the total keywords is then calculated. Since LLMs often extract appropriate keywords more effectively, kwPrec focuses more on factual relevance rather than expressional relevance. Fig. 3 illustrates the tokens segmented by our method compared to those obtained by BLEU-4 and ROUGE-L.
2.4 Automatic Labeling and Human Rechecking
Figure 4: Labeling and rechecking. (In Chinese: Fig. 12)
Automatic labeling
Human rechecking
We undertake a manual, one-to-one verification process by analyzing the annotated results and explanations provided by GPT4-0613 against the original news. This step ensures the accuracy of the machine-generated annotations. In the end, instances verified as accurate by annotators comprise the final $\mathbb{UHG}$Eval dataset. For details on manual annotation, please refer to Appendix A.1.
2.5 Dataset Statistics
Starting with 17,714 candidate hallucinated continuations, we curated a dataset of 5,141 hallucinated continuations, as detailed in the basic statistics in Table 2. For further analysis of statistics and an example of the dataset, please refer to the Appendix A.2 and the Appendix A.3, respectively.
DOC
KNO
NUM
GEN
#news
1242
320
2431
1148
avg. #hallu. kw.
2.15
1.99
2.54
2.12
avg. #kw.
8.43
8.09
8.07
8.17
#hallu. kw. / #kw.
25.47%
24.61%
31.44%
26.00%
avg. len. contn.
46.77
48.36
44.47
45.97
avg. len. begin.
102.15
102.66
103.20
102.86
avg. len. refer.
634.17
618.90
624.47
632.47
3 Experiments
3.1 Models
Given that our dataset is tailored for the Chinese language generation domain, we selected eight widely used Chinese LLMs and three models from OpenAI. These LLMs are from eight base models: Aquila2 Base BAAI (2023), Baichuan2 Base Yang et al. (2023a), GLM Base Du et al. (2022), GPT Base [^2] https://openai.com, InternLM Base InternLM (2023), Qwen Base Bai et al. (2023), BLOOMZ Base Muennighoff et al. (2023), and LLaMA2 Base Touvron et al. (2023). Refer to the Appendix B.1 for a detailed overview of the LLMs used in the experiments.
3.2 Evaluation Forms
In this study, we conducted a detailed analysis of evaluation methods across three dimensions: form, metric, and granularity. A more comprehensive report can be found in the Appendix B.2. Here, we introduce the three forms of evaluation.
3.3 Evaluation Framework
To accommodate different forms of evaluation methods, we have developed a data-secure, easy-to-extend, and easy-to-use evaluation framework, as illustrated in Fig. 5. Refer to Appendix B.3 for a more detailed understanding of the various layers of the framework.
Figure 5: Evaluation framework
3.4 Experimental Setup
To establish a robust experimental framework, we have set up some configurations as follows.
Prompt engineering
We apply the technique of “intent + instruction + 3-shot (explainable) prompting.” Intent delineates the role, instruction outlines the task, and the prompt incorporates three examples to aid the few-shot learning Zhao et al. (2023). Furthermore, political content in examples is prohibited to adhere to content policies from model service providers. Explainable prompting entails not merely acquiring results, but also eliciting the model’s rationale behind its responses. Refer to Appendix E to view the complete prompt templates.
Example Balancing
Hyperparameter settings
Metrics
3.5 Results and Analysis
Results are presented in Table 3 and Table 4.
Discriminative-Keyword Discriminative-Sentence Selective
avg. acc.
avg. #kws
#valid
avg. acc.
#valid
acc.
#valid
Aquila-34B
53.62%
3.00
3719
49.86%
5009
54.29%
4319
Baichuan2-13B
51.63%
3.128
4478
46.88%
5047
50.23%
5130
Baichuan2-53B
52.13%
2.98
1656
50.81%
1478
54.67%
4443
ChatGLM2-6B
50.80%
3.10
4289
43.87%
5130
43.59%
5130
GPT3.5-Turbo
53.72%
3.08
4183
50.02%
5039
49.03%
5103
GPT4-0613
70.04%
3.07
4100
57.42%
5024
55.20%
5047
GPT4-1106
[…]
3.10
4189
[…]
4903
60.35%
4752
InternLM-20B
50.92%
3.10
4388
51.01%
5130
49.43%
5130
Qwen-14B
52.86%
[…]
4478
50.58%
5130
54.74%
5130
Xinyu-7B
49.58%
3.12
4451
48.66%
5014
50.58%
5130
Xinyu2-70B
52.94%
3.12
4482
55.04%
5128
[…]
5129
Generative
avg. bleu
avg. rouge
avg. kwPrec
avg. bert
avg. len.
#valid
Aquila-34B
11.80%
6.04%
34.36%
67.51%
43.76
5130
Baichuan2-13B
8.84%
6.96%
25.51%
65.69%
46.04
5113
Baichuan2-53B
10.06%
[…]
26.45%
67.65%
49.40
3837
ChatGLM2-6B
9.17%
7.17%
24.53%
64.89%
46.27
5094
GPT3.5-Turbo
9.02%
6.30%
27.74%
66.39%
39.04
5084
GPT4-0613
10.74%
7.19%
28.47%
67.36%
44.41
5109
GPT4-1106
8.62%
6.86%
30.94%
67.38%
44.83
5121
InternLM-20B
14.89% 7.96%
31.10%
[…]
51.55
5125
Qwen-14B
12.72%
6.54%
32.95%
66.96%
45.85
5125
Xinyu-7B
10.30%
6.52%
28.64%
67.32%
49.84
4978
Xinyu2-70B
[…]
7.05%
[…]
68.97%
[…]
5130
KNO
DOC
GEN
NUM
Aquila-34B
59.55%
[…]
53.74%
53.52%
Baichuan2-13B
53.75%
[…]
48.43%
49.67%
Baichuan2-53B
57.70%
[…]
56.26%
52.58%
ChatGLM2-6B
40.94%
45.56%
[…]
42.63%
GPT3.5-Turbo
55.21%
[…]
47.63%
47.85%
GPT4-0613
59.87%
[…]
51.93%
55.73%
GPT4-1106
68.73%
60.19%
54.77%
[…]
InternLM-20B
51.88%
[…]
49.56%
48.43%
Qwen-14B
62.81%
[…]
53.15%
53.09%
Xinyu-7B
48.44%
52.02%
[…]
50.00%
Xinyu2-70B
63.13%
[…]
54.46%
57.07%
Discriminative evaluation
Selective evaluation
Firstly, GPT4-1106 clinches the top spot, reaffirming the formidable foundational capabilities of the GPT series models. Concurrently, Xinyu2-70B attains second place, excelling as a model trained on the Chinese news corpus. This achievement, to a degree, confirms the merit of domain-specific LLMs. Secondly, when comparing the outcomes of the selective evaluation with those of the discriminative evaluation at the sentence level, most LLMs exhibit improved accuracy. We think, furnishing LLMs with more contrasting information alleviates the demand for the model’s fact recall, thus diminishing the challenge of selective evaluation. Therefore, we posit that selective evaluation is comparatively simpler for LLMs.
Generative evaluation
Evaluations by Type
3.6 Further Discussion
@@ 4 Related Work
This section outlines benchmark datasets, their characteristics, and evaluation methodologies. These benchmarks are summarized in Table 5. For more related works, you may refer to Appendix C.
@@
KoLA-KC Yu et al. (2024) Direct: Wiki, evolving dataset Auto BLEU, ROUGE Document EN
@@ 4.3 Evaluation schemes
Currently, building automatic metrics for evaluation is still dominant, and a small proportion of works use human evaluation Min et al. (2023); Lin et al. (2022); Maynez et al. (2020). In terms of specific evaluation metrics, most works adopt common classification metrics, e.g., F1, accuracy, precision, and recall. some other works construct their calculation methods, e.g., FACTOR Muhlgay et al. (2023), FActScore Min et al. (2023), HaLoCheck Elaraby et al. (2023), etc. However, the above metrics are rule-based and can only evaluate the ability of LLMs to classify hallucinations, but not the ability of LLMs to generate content without hallucinations. Thus, some benchmarks explore even further in generative evaluation. For example, KoLA Yu et al. (2024) evaluates knowledge creation (KC) using BLEU and ROUGE, and TruthfulQA Lin et al. (2022) evaluates hallucinations using a specially trained classifier, GPT-judge.

### v3

@@ Abstract
Large language models (LLMs) produce hallucinated text, compromising their practical utility in professional contexts. To assess the reliability of LLMs, numerous initiatives have developed benchmark evaluations for hallucination phenomena. However, they often employ constrained generation techniques to produce the evaluation dataset due to cost and time limitations. For instance, this may involve employing directed hallucination induction or deliberately modifying authentic text to generate hallucinations. These are not congruent with the unrestricted text generation demanded by real-world applications. Furthermore, a well-established Chinese-language dataset dedicated to the evaluation of hallucinations is presently lacking. Consequently, we have developed an $\mathbb{U}$nconstrained $\mathbb{H}$allucination $\mathbb{G}$eneration Evaluation ($\mathbb{UHG}$Eval) benchmark, containing hallucinations generated by LLMs with minimal restrictions [^1] Framework, dataset, and results on our project webpage: https://iaar-shanghai.github.io/UHGEval/.. Concurrently, we have established a comprehensive benchmark evaluation framework to aid subsequent researchers in undertaking scalable and reproducible experiments. We have also evaluated prominent Chinese LLMs and the GPT series models to derive insights regarding hallucination.
@@ 1 Introduction
Large language models (LLMs) have unparalleled proficiency in language generation, knowledge application, and intricate reasoning Zhao et al. (2023). However, they invariably manifest hallucination Rawte et al. (2023); Yu et al. (2024c), as they often generate content that is incongruent with user input, the model’s output context, or factual information. Real-world hallucination examples from our $\mathbb{UHG}$Eval dataset can be observed in Fig. 1.
@@
While there have been a bunch of efforts to develop benchmarks for hallucination assessment, they always employ restricted techniques to produce particular kinds of hallucinated utterances. This approach is at odds with real-world scenarios where hallucinations arise in unrestricted, spontaneously generated content. For example, HaluEval specifies the type of hallucination in the prompt when generating hallucinated text: “You are trying to answer a question but misunderstand the question context and intention” Li et al. (2023). Additionally, benchmarks such as HaDes annotate hallucinations at a finer granularity by generating token-level hallucinations based on text perturbations Liu et al. (2022), but the text perturbation method is still constrained.
Hallucinations must be generated in an unconstrained setting; otherwise, it’s difficult to determine whether the hallucinated texts in many datasets are indeed errors that language models will make on their own. This point carries profound implications. For example, with a dataset containing freely generated hallucinations, researchers can explore the differences in model hidden states (logits, hidden layers, etc.) between hallucinated text spans and unhallucinated text spans. Such in-depth analysis would not be possible with datasets generated under constrained settings. Appendix A provides a detailed comparison with three other datasets, TruthfulQA Lin et al. (2022), HaluEval Li et al. (2023), and HaDes Liu et al. (2022).
Besides, many benchmarks are centered on the evaluation in English, neglecting the assessment of hallucination in Chinese. The extensive lexicon of Chinese characters, combined with the complexities introduced by Chinese word segmentation, renders the Chinese hallucination evaluation particularly arduous and deserving of focused scrutiny.
@@
Our contributions: (1) The development of an unconstrained hallucination evaluation dataset, comprising over 5000 items. Existing methods for constructing datasets often yield biases towards predefined directions, thereby hindering the full simulation of real-world hallucinations. (2) The establishment of a unified and diverse evaluation framework, $\mathbb{UHG}$Eval, that encompasses discriminative, selective, and generative evaluations. Current benchmark methods for hallucination evaluation often exhibit a singular approach and lack task specificity. (3) A comprehensive empirical analysis. We evaluated eight prominent Chinese LLMs and three classic GPT series models to explore the credibility of various LLMs.
@@ 2 Related Work
This section outlines hallucination evaluation benchmarks, their characteristics, and evaluation methodologies. A summary of these benchmarks is presented in Table 1. For related works on LLMs and hallucinations, please refer to Appendix B.
@@
KoLA-KC Yu et al. (2024a) Direct: Wiki, evolving dataset Auto BLEU, ROUGE Document EN
@@ 2.3 Evaluation Schemes
Currently, building automatic metrics for evaluation is still dominant, and a small proportion of works use human evaluation Min et al. (2023); Lin et al. (2022); Maynez et al. (2020). In terms of specific evaluation metrics, most works adopt common classification metrics, e.g., F1, accuracy, precision, and recall. Some other works construct their calculation methods, e.g., FACTOR Muhlgay et al. (2024), FActScore Min et al. (2023), HaLoCheck Elaraby et al. (2023), etc. However, the above metrics are rule-based and can only evaluate the ability of LLMs to classify hallucinations, but not the ability of LLMs to generate content without hallucinations. Thus, some benchmarks explore further in generative evaluation. For example, KoLA Yu et al. (2024a) evaluates knowledge creation (KC) using BLEU and ROUGE, and TruthfulQA Lin et al. (2022) evaluates hallucinations using a specially trained classifier, GPT-judge.
3 The UHGEval Dataset
3.1 Data Collection and Pre-processing
Type Share Categories
DOC 27.52% Politics, Law, Military, Education
NUM 43.34% Sports, Economy, Market
KNO 6.55% Science, Technology, Healthcare
GEN 22.59% Society, Culture, Arts, Entertainment, Weather, Environmental Protection, Disasters, Accidents
3.2 Unconstrained Hallucination Generation
Furthermore, current benchmarks for evaluating hallucination have predominantly relied on a single LLM to produce a hallucinated dataset. Notable examples include HaluEval Li et al. (2023) and PHD Yang et al. (2023b), which exclusively utilize ChatGPT, and FActScore Min et al. (2023) and FACTOR Muhlgay et al. (2024), which solely employ InstructGPT Ouyang et al. (2022). In contrast, our methodology incorporates a suite of five distinct Chinese LLMs to generate hallucinated content. These models include ChatGLM2-6B Du et al. (2022), Baichuan2-13B Yang et al. (2023a), Qwen-14B Bai et al. (2023), InternLM-20B InternLM (2023), and Xinyu-7B. For additional information about the Xinyu series models, please refer to the Appendix D.1.
For each input news article, we concurrently generate five candidate continuations using five different LLMs without constraint. Overall, our approach engenders a more unconstrained and heterogeneous generation of hallucinations, mitigating the bias that may arise from the use of a single model or constrained prompting.
3.3 Hallucination Ranking
Hallucination ranking is a crucial step in selecting the most appropriate continuation from the five candidates generated by the five LLMs. This process relies on two critical metrics: $fluency$, ensuring that the continuation does not become too nonsensical, and $likelihood$, which stands for the likelihood of hallucination occurrence, ensuring that the continuation includes a detectable level of hallucinations. They are computed as follows.
Fluency
This refers to the coherence and readability of the text Liang et al. (2024). A fluent text should read smoothly and be grammatically correct in the context of the continuation. To assess fluency, a reward model developed by the Institute for Advanced Algorithms Research (IAAR) is employed, trained to score text fluency. The model is fine-tuned using a dataset annotated with news on an open-source reward model, Ziya model [^2] https://huggingface.co/IDEA-CCNL/Ziya-LLaMA-7B-Reward.
Likelihood of Hallucination Occurrence
kwPrec uses an LLM (e.g., GPT3.5-Turbo) to extract keywords from the continuation and then determine whether these keywords have exact matches in the reference information. The ratio of all matches to the total keywords is then calculated. Since LLMs often extract appropriate keywords more effectively, kwPrec focuses more on factual relevance rather than expressional relevance. Fig. 3 illustrates the tokens segmented by kwPrec compared to those obtained by BLEU-4 and ROUGE-L. The prompt template utilized for extracting keywords is depicted in Fig. 13 within Appendix F.
Algorithm 1 Hallucination Ranking
$candidate:list[str]$
$final:str$
$candidate.sort(descend,by=fluency)$
$picked\leftarrow candidate[:3]$triangleright More fluency
$picked.sort(ascend,by=kwPrec)$
$final\leftarrow picked[0]$triangleright More Hallucination
With $fluency$ and $kwPrec$, our task in the hallucination ranking step is to select one out of five candidate continuations that appears to be correct (highest in $fluency$) but is likely to contain hallucinations (lowest in $kwPrec$).
The specific steps are as follows (also shown in Algorithm 1). Step 1: Rank the five candidate continuations in descending order by $fluency$. Step 2: Select the top three continuations with the highest $fluency$. Step 3: Rank these three continuations in ascending order by $kwPrec$. Step 4: Choose the continuation with the lowest $kwPrec$ score. Following these steps, the continuation selected in Step 4 is the $final$ choice. By employing such a ranking, it is guaranteed that, in the worst-case scenario, the $final$ candidate ranks at least third in fluency and third in the likelihood of hallucination occurrence, achieving a balanced level.
3.4 Automatic Labeling and Human Rechecking
Figure 4: Labeling and rechecking. (In Chinese: Fig. 12)
Automatic labeling
Human rechecking
We undertake a manual, one-to-one verification process by analyzing the annotated results and explanations provided by GPT4-0613 against the original news. This step ensures the accuracy of the machine-generated annotations. In the end, instances verified as accurate by annotators comprise the final $\mathbb{UHG}$Eval dataset. For details on manual annotation, refer to Appendix C.1.
3.5 Dataset Statistics
Starting with 17,714 candidate hallucinated continuations, we curated a dataset of 5,141 hallucinated continuations, as detailed in the basic statistics in Table 3. For further analysis, the data volume of each step in the dataset creation pipeline, and an example of the dataset, please refer to Appendix C.2, Appendix C.3 and Appendix C.4, respectively.
DOC KNO NUM GEN
#news 1242 320 2431 1148
avg. #hallu. kw. 2.15 1.99 2.54 2.12
avg. #kw. 8.43 8.09 8.07 8.17
#hallu. kw. / #kw. 25.47% 24.61% 31.44% 26.00%
avg. len. contn. 46.77 48.36 44.47 45.97
avg. len. begin. 102.15 102.66 103.20 102.86
avg. len. refer. 634.17 618.90 624.47 632.47
4 Experiments
4.1 Models
Given that our dataset is tailored for the Chinese language generation domain, we selected eight widely used Chinese LLMs and three LLMs from OpenAI. These LLMs are from eight base models: Aquila2 BAAI (2023), Baichuan2 Yang et al. (2023a), GLM Du et al. (2022), GPT [^3] https://openai.com, InternLM InternLM (2023), Qwen Bai et al. (2023), BLOOMZ Muennighoff et al. (2023), and LLaMA2 Touvron et al. (2023). Refer to the Appendix D.1 for a detailed overview of the LLMs used in the experiments.
4.2 Evaluation Forms
In this study, we conducted a detailed analysis of evaluation methods across three dimensions: form, metric, and granularity. A more comprehensive report can be found in the Appendix D.2. Here, we introduce the three forms of evaluation.
4.3 Evaluation Framework
To accommodate different forms of evaluation methods, we have developed a data-secure, easy-to-extend, and easy-to-use evaluation framework, as illustrated in Fig. 5. Refer to Appendix D.3 for a more detailed understanding of the various layers of the framework.
Figure 5: Evaluation framework
4.4 Experimental Setup
Prompt Engineering
We apply the technique of “intent + instruction + 3-shot (explainable) prompting.” Intent delineates the role, instruction outlines the task, and the prompt incorporates three examples to aid the few-shot learning Chen et al. (2024); Yu et al. (2024b). Furthermore, political content in examples is prohibited to adhere to content policies from model providers. Explainable prompting entails not merely acquiring results, but also eliciting the model’s rationale behind its responses. Refer to Appendix F to view the complete templates.
Example Balancing
Hyperparameter Settings
Metrics
4.5 Results and Analysis
Results are presented in Table 4 and Table 5.
Discriminative-Keyword Discriminative-Sentence Selective
avg. acc. avg. #kws #valid avg. acc. #valid acc. #valid
Aquila-34B 53.62% 3.00 3719 49.86% 5009 54.29% 4319
Baichuan2-13B 51.63% 3.128 4478 46.88% 5047 50.23% 5130
Baichuan2-53B 52.13% 2.98 1656 50.81% 1478 54.67% 4443
ChatGLM2-6B 50.80% 3.10 4289 43.87% 5130 43.59% 5130
GPT3.5-Turbo 53.72% 3.08 4183 50.02% 5039 49.03% 5103
GPT4-0613 70.04% 3.07 4100 57.42% 5024 55.20% 5047
[…]
InternLM-20B 50.92% 3.10 4388 51.01% 5130 49.43% 5130
[…]
Xinyu-7B 49.58% 3.12 4451 48.66% 5014 50.58% 5130
[…]
Generative
avg. bleu avg. rouge avg. kwPrec avg. bert avg. len. #valid
Aquila-34B 11.80% 6.04% 34.36% 67.51% 43.76 5130
Baichuan2-13B 8.84% 6.96% 25.51% 65.69% 46.04 5113
[…]
ChatGLM2-6B 9.17% 7.17% 24.53% 64.89% 46.27 5094
GPT3.5-Turbo 9.02% 6.30% 27.74% 66.39% 39.04 5084
GPT4-0613 10.74% 7.19% 28.47% 67.36% 44.41 5109
GPT4-1106 8.62% 6.86% 30.94% 67.38% 44.83 5121
[…]
Qwen-14B 12.72% 6.54% 32.95% 66.96% 45.85 5125
Xinyu-7B 10.30% 6.52% 28.64% 67.32% 49.84 4978
[…]
KNO DOC GEN NUM
[…]
Discriminative Evaluation
Selective Evaluation
Firstly, GPT4-1106 clinches the top spot, reaffirming the formidable foundational capabilities of the GPT series models. Concurrently, Xinyu2-70B attains second place, excelling as a model trained on the Chinese news corpus. This achievement, to a degree, confirms the merit of domain-specific LLMs. Secondly, when comparing the outcomes of the selective evaluation with those of the discriminative evaluation at the sentence level, most LLMs exhibit improved accuracy. We think, furnishing LLMs with more contrasting information alleviates the demand for the model’s fact recall, thus diminishing the challenge of selective evaluation. Therefore, we posit that selective evaluation is comparatively simpler for LLMs. Thirdly, a decline is observed in discriminative evaluation outcomes from GPT4-0613 to GPT4-1106, whereas selective evaluation outcomes register a notable increase of around 5%. This substantiates the “seesaw phenomenon,” wherein certain capabilities are enhanced while others may regress, in tandem with the model’s upgrade Zheng et al. (2023). This suggests that the decision to either enhance a single capability individually or to balance multiple capabilities is critical.
Generative Evaluation
Evaluations by Type
4.6 Further Discussion
Moreover, various evaluations cater to different application contexts. Should the objective be to solely improve the model’s capacity for reliable continuation, generative evaluation would suffice. In the training of a dependable chatbot, selective and discriminative evaluations prove suitable. When aiming to train a reward model, selective evaluation is beneficial, offering evaluation for positive and negative instances. If the goal is to enhance the model’s ability to recall and apply knowledge, discriminative evaluation emerges as the demanding option.